6:30PM: Sky condition: Cirrus around, reasonably clear wind: ~10m/s humidity: 41%

Calibration taken for two masks for UC 1795.25 and IC1795.26 (0002-0019)

Collimation problem since ~6:40PM..........

7:20pm: seeing >2.5" but improving to 1.6-1.8" seeing 7:25pm: slewing to target field (first field: mods.26.IC1795)

modsAlign 5 alignment stars are little off from center of the boxes, but visible at edges. go ahead with obs.

3 of the 7 exposures obtained with >1.5" seeing and collimation problem (getting worse). (image names: 0020-0023) perhaps usable for bright sources .

8:45pm: seeing 1.45" guide image is still not good... trying to improve

8:50pm: after 3rd exposure we aborted to improve collimation. WFS view showed image that looked like triangle. (at 10PM when we resumed it looked better.)

9:25pm: still struggling with collimation. wind: ~10m/s, humidity 18%

9:45pm: Discussing problems with the collimation issue. dome seeing? why? secondary won't flatten (doug) John Hill has other idea?

10pm: stil working on collimation. may be secondary problem?

10:10pm: starting mods.38 (IC1795 F2). try 5 exposures???

10:15pm: science field mods.38 starts. avg. seeing: 1.5" cirrus around whole half a night.

Problem is not with MODS but with telescope... Secondary got into some kind of wacky stage; had coma, and were trying to correct for coma? Perhaps combination with bad seeing.

10:50pm: temperature rising slowly. now 2.3c wind 11-15m/s, humidity 13%, seeing 1.1" (getting better) 11:05pm: seeing 1.3", wind: 7-15m/s, temp:2.6c, humidity: 13.3 11:20pm: seeing 1.4" 12:05am: Kim et al. exposure done after 5th exposure of IC1795 F2.

seeing: 1.5"

Comments: Guiding needs to be improved... This needs to be in high priority.

written by Serena Kim

================================= J Mauerhan takes over at 12:15

12:35am: seeing 0.8" 4:20 am: seeing degraded to 1.1-1.3"

5:15am collimation problems after moving to lower elevation source (50 deg elevation). required manual intervention by doug, but not fully resolved by end of night.

was able to get good data on 3 SNe.

-- XiaohuiFan - 26 Jan 2012
Topic revision: r4 - 26 Jan 2012, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback